OpenStudy (anonymous):

Derek kicks a soccer ball off the ground and in the air, with an initial velocity of 31 feet per second. Using the formula H(t) = -16t^2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anwaarullah):

at max height H(t)=0 -32t+v=0 t=v/32 t=31/32 H(31/32)=-16(31/32)^2+31*31/32 =15 feet